# King’s College London, U.K

### v4

@@ A.3.1 Assembly index defines an encoding scheme
    Let $\Gamma=\langle V_{\Gamma},E_{\Gamma},s_{\gamma},t_{\gamma}\rangle$ be the acyclic finite quiver with basic objects $\Sigma_{G}\subset V_{\Gamma}$ such that was defined to compute the assembly index $c_{\Gamma}(x)$. There exists a CFG $G=\langle V_{G},\Sigma_{G},R,S\rangle$ such that $V_{\Gamma}-\Sigma=V_{G}$, $\Sigma_{\Gamma}=\Sigma_{G}$ and $|E_{\Gamma}|=|R|$.
@@
    Let $\Gamma$ be the acyclic finite quiver that encodes the generation of the string $x$. First, we define the set of terminal symbols in $G$ as the set of basic objects in $\Gamma$, $\Sigma_{G}=\Sigma_{\Gamma}=\Sigma$. Let $V_{G}=(V_{\Gamma}-\Sigma)\cup\{S\}$ be the set of non-terminal symbols.
@@
    Now, we construct $R$ in the following way: We start with $S\rightarrow x$. Now, for each $e_{i}\in E$ such that $s_{\gamma}(e_{i})=x_{i}$ and $s_{\gamma}(e_{i})=y_{i}$, we define the rule as $y_{i}\rightarrow x_{i}z_{i}$ or $y_{i}\rightarrow x_{i}z_{i}$, $z_{i}\in V_{\Gamma}$, whichever concatenation of combinations builds the object encoded by vertex $y_{i}$. This is possible since each “assembly step” can only combine two objects. If the combination was unrestricted, then the assembly index of strings would always be one, as you can combine all the characters in the first step. For any other limit we can use a rule with more symbols.
    It is easy to prove by induction over the number of vertices $|V_{\Gamma}|$ that $G$ produces $x$. If $|V_{\Gamma}|=1$, then $x$ must be basic in $\Gamma$, therefore terminal in $G$. It follows that $S\rightarrow x$ generates $x$. Now we assume that the theorem holds for $n$ or less; consider the quiver $V_{\Gamma}={x_{1},\dots,x_{n},x}$ so that $\Gamma$ generates $x$. By hypothesis there exists an edge that, for some $x_{i}$, goes to $x$. Therefore there is the rule $x\rightarrow x_{i}z_{i}$ or $X\rightarrow z_{i}x_{i}$ such that $x$ is either concatenation. By hypothesis, the subgrammars defined by removing $x$ and adding the rule $S\rightarrow z_{i}$ and $S\rightarrow x_{i}$ must generate $z_{i}$ and $x_{i}$. Therefore we have the rules to generate $z_{i}$ and $x_{i}$; thus with the rule $x\rightarrow x_{i}z_{i}$ or $x\rightarrow z_{i}x_{i}$ we can generate $x$. Then, since $S\rightarrow X$, $\Gamma$ generates $x$.
    Finally, we remove the rule $S\rightarrow x$ and replace $x$ by $S$, which is equivalent to just defining the starting symbol as being $x$. By construction, the stated equities between the sets hold.
@@
    The grammar defined in Theorem 7 demonstrates that any representation of a string’s assembly by an acyclic finite directed graph can be encoded into a Context-Free Grammar (CFG) in Chomsky Normal Form (CNF), capturing the same information regarding the relationship of how a string is constructed from basic components. The assembly index is defined as the number of non-basic vertices in the graph, while the size of the grammar is defined as the number of rules, which is equivalent to the number of nonterminals if there are no useless rules [55]. Let’s denote this size by $|G|$. Cronin et al. define the assembly space as a quiver. It is not clear that there is any necessity for having two or more edges linking the same pair of vertices in an assembly space, but for completeness’ sake, we will denote by $G^{*}$ the grammar with any useless rules removed. We will show that there is a simple linear relationship between the assembly index $c_{\Gamma}(x)$ and the size of the grammar $|G^{*}|$.
@@
    Let $x$ be a string with minimal assembly space $\Gamma$ and $G$ be the CFG in Theorem 7. Then
    • $c_{\Gamma}(x)=|G^{*}|$.
[…]
@@
    The first equation is by construction of $G$. The second is a direct consequence of the first and [55, Theorem $1$]. ∎
    With the stated corollary we have proven that the assembly index for strings is equivalent to the compression length of a compressing grammar and, in the optimal case it can only converge towards LZ compression [30, 59, 60] and thus is bounded by Shannon Entropy by the (noiseless) Source Coding Theorem [61, 58]. It is important to note that the inequality arises when the grammar, thus the pathway assembly, is not optimal at identifying the ”building blocks” of $x$ given the same alphabet of basic objects. For compressing grammars in particular, several implementation-specific entropy bounds were found in [56].
@@ A.3.3 Assembly index is a version of LZ compression
    In Section A.3, we have demonstrated that either the set of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ or at least one of the longest rooted paths in this ${\Gamma^{*}}_{y}$ is sufficient to encode the object $y$. These results employ general algorithmic methods so that they are independent of the choice of the programming language and encoding-decoding schemes. In Theorem 7, we have demonstrated that every finite assembly space can be generated by a corresponding context-free grammar (CFG), which reveals a way one can compress an assembly space (e.g., the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ that generates the object $y$) into the minimum set of rules of this CFG. These results together reveal the existence of some LZ-based method to encode the object and its assembling process.
    Indeed, the proof of the following Theorem 9 constructs such a LZ compression scheme. In practice, the major characteristic that distinguishes assembly index compression from other traditional LZ schemes, is that: in the latter case, one would parse the string into contiguous smaller strings (or codewords) whose resulting unidimensional concatenation directly represents the string; while in the former case, one would parse the object into a non-unidimensional structure where each block is “concatenated” according to a directed acyclic multigraph. Thus, as shown in Theorem 9, a LZ compression scheme from the assembly index requires a way to compress this non-linear structure (in particular, a directed acyclic multigraph) back into a linear sequence of codewords.
    For LZ-based schemes, one has a sequence of encoded tuples of finite length. Typically, each one of these tuples contains at least one pointer (e.g., in the case of LZ78787878 and LZW, called indexes) that represents to another tuple. Also, it contains information about some elements of the basic alphabet (e.g., in the case of LZ78787878 and LZW, it contains the codeword for only one symbol to be concatenated at the end of the element the pointer is referring to). In the same fashion, one expects that the sequence of edges that constitutes an assembling path (from the basis set $B_{\Gamma}$ to the terminal object/vertex in a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$) can be converted into a sequence of respective encoded tuples that contains pointers, codewords for other objects, etc. One can think of these assembling paths being, for example, the sequence defined by either a longest rooted path $\gamma_{x}^{\max}$ (as investigated in Section A.3.1) or a shortest rooted path $\gamma_{x}^{\min}$, where both belong to the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$.
@@
    The pointers in LZ compression schemes traditionally refer to encoded tuples that have occurred before. For Assembly Theory, the pointers may refer to other tuples that have not occurred before in an assembling path, but are part of the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$, that is, part of a “tree”-like structure. We demonstrate in Theorem 9 that the minimality of ${\Gamma^{*}}_{x}$ in fact guarantees that such a scheme, where pointers may not refer to the unidimensional past subsequences, allows unique decoding, despite the non-unidimensional structure of assembly spaces.
    Theorem 9 (LZ encoding from the assembly index).
@@
    Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly index of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $\Gamma\subset\mathcal{S}$ with $y\in V\left(\Gamma\right)$, $\gamma_{y}^{\min}$, and $\gamma_{y}^{\max}$ be arbitrary. Then, one has it that the sequence $S_{AT}\left(y\right)$ composed of $\left|\gamma_{y}^{\min}\right|$ encoded tuples in the form $c,i_{t},v_{l}\right>$ (except for the 00-th initial tuples which are $t\right>$, $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, and $v_{B}\right>$, respectively, where $v_{B}\in B_{\Gamma}$ and $t\in\mathbb{P}\left(\mathcal{T}\right)$) encodes the object $y$ such that
@@
    • $t\right>$ is an encoding of $t\in\mathbb{P}\left(\mathcal{T}\right)$, where $\mathbb{P}\left(\mathcal{T}\right)$ is the powerset of possibles edges of the minimum directed acyclic multigraph $\mathcal{T}$ into which any possible minimum rooted assembly subspace whose longest rooted path has length $\left|\gamma_{y}^{\max}\right|$ can be embedded, and all the most distant vertices from the source can only be the vertices in $B_{\Gamma}$.
@@
    • $v_{l}$ is another index/pointer that refers to the edge $i_{t}$ whose source vertex $s\left(i_{t}\right)=v\in V\left(\Gamma\right)$ is the vertex that is the label of the edge $i_{t}$ when $\Gamma$ is embedded into $\mathcal{T}$;
@@
    • $\;c_{\Gamma}\colon\Gamma\subset\mathcal{S}\to\mathbb{N}\;$ gives the assembly index of the object $y$ in the assembly space $\Gamma$ (or $\mathcal{S}$);
@@
    as input. Then, it constructs the directed acyclic multigraph (DAmG) $\mathcal{T}$ of height $\left|\gamma_{y}^{\max}\right|$, and enumerates all the possible minimum rooted assembly subspaces contained in $\mathcal{T}$: starting from those that are isomorphic to $\gamma_{y}^{\max}$ with any possible combination of initial vertex/object in $B_{\Gamma}$; and then in a dovetailing procedure, indexing first those subspaces that differ (from $\gamma_{y}^{\max}$) by the least amount of additional edges. In the next step, program $p_{5}$ reads the rest of the sequence $S_{AT}\left(y\right)$, extracting from the encoded tuples the set [^2] Note that some of these members may actually refer to the same index. Thus, the size of this set may be much smaller than $2\left|\gamma_{y}^{\min}\right|$. $I=\left\{{i_{t}}_{1},\dots,{i_{t}}_{\left|\gamma_{y}^{\min}\right|},{v_{l}}_{1},\dots,{v_{l}}_{\left|\gamma_{y}^{\min}\right|}\right\}$ of indexes of edges that are contained in $\mathcal{T}$. Then, by employing $\mathbf{F}$, program $p_{4}$ enumerates all possible minimum assembly subspaces embedded into $\mathcal{T}$ that necessarily contain every edge in $I$, thereafter choosing the $t$-th first $\Gamma^{\prime}$ that appears in this enumeration. Finally, it uses $\Gamma^{\prime}$ as input to program $p_{3}$ from the proof of Theorem 5 to return the final object $y^{\prime}$. This describes the decoding procedure from $S_{AT}\left(y\right)$. In order to construct a program $p^{\prime}_{5}$ that encodes a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ into a $S_{AT}\left(y\right)$, one just needs to note that the codewords for $t\right>$, $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, $v_{B}\right>$, $\gamma_{y}^{\min}\right>$, and $y\right>$ are always retrievable by an algorithm given ${\Gamma^{*}}_{y}$ as input. From these, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to construct $\mathcal{T}$ and enumerate its possible minimum rooted assembly subspaces. Then, it extracts $I$ from $\mathcal{T}$ by matching those that correspond to the edges and their respective labels in $\gamma_{y}^{\min}$. Next, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to enumerate the minimum rooted assembly subspaces embedded into $\mathcal{T}$ which contains every edge in $I$ so as to find the index $t$ that corresponds to ${\Gamma^{*}}_{y}$ in this enumeration. Finally, the program returns the sequence $S_{AT}\left(y\right)$ as output. ∎
@@
    With the purpose of demonstrating that $\Gamma^{\prime}={\Gamma^{*}}_{y}$, one can exploit the minimality of ${\Gamma^{*}}_{y}$ and $\gamma_{y}^{\min}$ and the maximality of $\gamma_{y}^{\max}$. First, note that there is only a finite number of rooted assembly spaces whose most distant vertices belong to $B_{\Gamma}$, which is also finite and fixed given $\Gamma$, and are separated by $\left|\gamma_{y}^{\max}\right|$ hops. From any rooted assembly space one of whose longest rooted paths is $\gamma_{y}^{\max}$, in order to construct $\mathcal{T}$, an algorithm can minimise the number of sufficient and necessary new vertices that needs to be added to $\gamma_{y}^{\max}$ by the following procedure: assign an index $j\in\mathbb{N}$ to each of the edges in $\gamma_{y}^{\max}$ by starting from the closest ones to the farthest ones from the initial vertex, which is the object in the basis set $B_{\Gamma}$, such that $1\leq j\leq\left|\gamma_{y}^{\max}\right|$; for the label of the edge $j=1$ in $\gamma_{y}^{\max}$, one only requires at most one extra vertex, and this vertex can only belong to $B_{\Gamma}$ because otherwise, since there would need to be an additional edge connecting this new vertex to the target vertex of $j=1$, it would contradict the maximality of $\gamma_{y}^{\max}$; the same analogously holds for each arbitrary $j+1\leq\left|\gamma_{y}^{\max}\right|$, so that labelling this edge $j+1$ with a vertex only requires at most one extra new vertex, and this vertex can only be at level $\leq j$, i.e., it can only belong to the set of vertices that are separated from the terminal vertex $y$ by at least $j+1$ hops; and for each new vertex created at level $j$ one would only need at most $2$ new vertices at level $\leq j-1$, and so on. Thus, by induction, one will have it that

### v5

@@ A.3.1 Assembly index defines an encoding scheme
Let ${\Gamma^{*}}_{x}$ be an arbitrary minimum rooted assembly (sub)space from which the assembly index $c_{\Gamma}(x)$ is calculated. Then, there exists a CFG $G=\left(V_{G},\Sigma_{G},R,S\right)$ that generates $x$ such that $V\left({\Gamma^{*}}_{x}\right)\setminus\left\{x\right\}=V_{G}\setminus\Sigma_{G}$, $\Sigma_{G}=B_{\Gamma}\cup\left\{x\right\}$, and
$\left|E\left({\Gamma^{*}}_{x}\right)\right|+\left|B_{\Gamma}\right|=|R|$ (10)
hold, where $B_{\Gamma}$ is the basis set of the assembly space $\left(\Gamma,\phi\right)$ from which $x$ is one of its vertices/objects such that ${\Gamma^{*}}_{x}\subseteq\Gamma$.
@@
Let ${\Gamma^{*}}_{x}$ generates $x$. First, we define the set of terminal symbols in $G$ as the set of basis objects in ${\Gamma^{*}}_{x}$ plus the final object $x$ to be assembled, that is, $\Sigma_{G}=B_{\Gamma}\cup\left\{x\right\}$. Then, let $V_{G}\setminus\Sigma_{G}=V\left({\Gamma^{*}}_{x}\right)\setminus\left\{x\right\}$ be the set of non-terminal symbols. Now, we construct the set $R$ of derivation rules of the CFG in the following way: we start by adding $S\rightarrow x$ for each basis object $x$ that belongs to $V_{G}\setminus\Sigma_{G}$ and also for each basis object that belongs to $\Sigma_{G}$ to the set of derivation rules666 Thus, there will be $2\left|B_{\Gamma}\right|$ rules in the form $S\rightarrow x$.; then, for each $e_{i}\in E\left({\Gamma^{*}}_{x}\right)$ such that $s_{\Gamma}(e_{i})=x_{i}$ and $t_{\Gamma}(e_{i})=y_{i}$, we construct a corresponding new rule, either $x_{i}\rightarrow x_{i}z_{i}$ or $x_{i}\rightarrow z_{i}x_{i}$, where $z_{i}\in V\left({\Gamma^{*}}_{x}\right)$ and $\phi\left(e_{i}\right)=z_{i}$, whichever concatenation of combinations builds the object/vertex $y_{i}$ (i.e., either $y_{i}=x_{i}z_{i}$ or $y_{i}=z_{i}x_{i}$). This is possible since each “assembly step” (represented by the respective edge) can only combine two objects: one object is the source vertex of the edge; and the other object is the vertex that is the label of this edge. If $y_{i}=x_{i}z_{i}\neq x$ (or $y_{i}=z_{i}x_{i}\neq x$), then we add $y_{i}$ to the set of non-terminal symbols $V_{G}\setminus\Sigma_{G}$. Otherwise, we add $y_{i}=x$ to the set of terminal symbols of $G$. From this construction of the set $R$, one will obtain that
$\left|E\left({\Gamma^{*}}_{x}\right)\right|+\left|B_{\Gamma}\right|=|R|\text{ .}$ (11)
The rest of the proof that $G$ generates $x$ follows by induction over the number of vertices in $V\left({\Gamma^{*}}_{x}\right)$. If $\left|V\left({\Gamma^{*}}_{x}\right)\right|=1$, then $x$ must be a basis object in $B_{\Gamma}$. Hence, $x$ is a terminal symbol in $G$, and thereby it directly follows that $S\rightarrow x$ generates $x$. Now, we assume that $G$ generates any non-terminal symbol $x_{i}$ for $i\leq n$ with a set $R_{n}$ of derivation rules, where each vertex $x_{i}$ belongs to a rooted path (in ${\Gamma^{*}}_{x}$) whose terminal vertex is $x$. Let $V\left({\Gamma^{*}}_{x}\right)=\left\{x_{1},\dots,x_{n},x_{n+1},\dots,x\right\}$ be the set of vertices so that ${\Gamma^{*}}_{x}$ generates $x$. By the inductive hypothesis, since there exists an edge $e$ with $s_{\Gamma}\left(e\right)=x_{i}$ and $t_{\Gamma}\left(e\right)=x_{j}$ for some $x_{i}$ and $x_{j}$ with $i,j\leq n$, we have it that there is the rule $x_{i}\rightarrow x_{i}z_{i}$ or $x_{i}\rightarrow z_{i}x_{i}$, where either $x_{j}=x_{i}z_{i}$ or $x_{j}=z_{i}x_{i}$, respectively. By construction of ${\Gamma^{*}}_{x}$, we have it that there is a vertex $x^{\prime}_{i}\in\left\{x_{1},\dots,x_{n}\right\}$ that is the source vertex of an edge $e^{\prime}\in E\left({\Gamma^{*}}_{x}\right)$ whose label is another $x^{\prime}_{j}\in\left\{x_{1},\dots,x_{n}\right\}$ and the target vertex of $e^{\prime}$ is $t_{\Gamma}\left(e^{\prime}\right)=x_{n+1}$. Then, we add the rule $x^{\prime}_{i}\rightarrow x^{\prime}_{i}x^{\prime}_{j}$ or $x^{\prime}_{i}\rightarrow x^{\prime}_{j}x^{\prime}_{i}$ to $R_{n}$, where $x_{n+1}=x^{\prime}_{i}x^{\prime}_{j}$ or $x_{n+1}=x^{\prime}_{j}x^{\prime}_{i}$, respectively. Therefore, since by the inductive hypothesis both $x^{\prime}_{i}$ and $x^{\prime}_{j}$ can be generated by the set $R_{n}$ of rules, we will have it that $R_{n}\cup\left\{x^{\prime}_{i}\rightarrow x^{\prime}_{i}x^{\prime}_{j}\lor x^{\prime}_{i}\rightarrow x^{\prime}_{j}x^{\prime}_{i}\right\}$ generates the object $x_{n+1}$. Finally, by the above construction of the set $R$ of derivation rules, we will have it that if $x_{n+1}=x$, then $x_{n+1}$ will be a terminal symbol and $G$ stops exactly at the object $x$.
@@
The grammar defined in Theorem 7 demonstrates that any assembled object can be encoded into a Context-Free Grammar (CFG) in Chomsky Normal Form (CNF) that captures the same information regarding the relationship of how an object is constructed from other objects. The assembly index is defined as the number of non-basis vertices, while the size of the grammar is defined as the number of derivation rules, which is equivalent to the number of nonterminals if there are no useless rules [15]. Let’s denote this size by $|G|$. In order to encode the object $x$ assembled according to ${\Gamma^{*}}_{x}$, there is no necessity for having two or more edges linking the same pair of vertices in an assembly space. Thus, for completeness’ sake, we will denote by $G^{*}$ the grammar with any useless rules/edges removed. We will show in Corollary 8 that there is a simple linear relationship between the assembly index $c_{\Gamma}(x)$ and the size of the grammar $|G^{*}|$, except for the size of the basis set $B_{\Gamma}$ which remains constant for any object assembled upon combinations of basis objects.
@@
Let $x$ be a string/object assembled according to ${\Gamma^{*}}_{x}$ and $G$ be the CFG in Theorem 7. Then,
  • the assembly index is equivalent to the non-basis symbols in the minimum grammar that generates $x$, i.e.,
    $c_{\Gamma}(x)=|G^{*}|-\left|B_{\Gamma}\right|\text{ ;}$ (12)
  • the number of factors in the LZ encoded form of $x$, denoted by $\text{LZ}(x)$, is upper bounded by the assembly index (except for the basis set $B_{\Gamma}$ that is finite and constant), i.e.,
    $LZ(x)\leq c_{\Gamma}(x)+\left|B_{\Gamma}\right|\text{ .}$ (13)
@@
The first equation is by construction of $G$ in Theorem 7. The second is a direct consequence of the first and [15, Theorem $1$]. ∎
Corollary 8 proves that the assembly index for strings is equivalent to the compression size of a compressing grammar, demonstrating that the notion of the minimum “amount of assembling steps” (that the minimum rooted assembly subspace defines) is exactly the same as a (minimal) CFG. For each of the grammars constructed in Theorem 7, the final object $x$ is unique, and for each distinct assembling path leading to $x$, there is a (unique) correspondent sequence of rules in $R$ that also results in the object $x$.
In the optimal case, such a size can only converge towards a traditional LZ compression scheme [35, 59, 60], such as LZ77777777 and LZ78787878, and thus its compression rate capabilities are bounded by Shannon Entropy by the (noiseless) Source Coding Theorem [61, 58]. It is important to note that the strict inequality arises when the grammar, thus the pathway assembly, is not optimal at identifying the “building blocks” of $x$ given the same alphabet of basis objects. For compressing grammars in particular, several implementation-specific entropy bounds were found in [56].
@@ A.3.3 Assembly index is a version of LZ compression
In Section A.3, we have demonstrated that either the set of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ or at least one of the longest (or shortest) rooted paths in this ${\Gamma^{*}}_{y}$ is sufficient to encode the object $y$. These results employ general algorithmic methods so that they are independent of the choice of the programming language and encoding-decoding schemes. In Theorem 7, we have demonstrated that every minimum rooted assembly (sub)space corresponds to at least one context-free grammar (CFG), which reveals a way one can compress an assembly space (e.g., the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ that generates the object $y$) into the minimum set of rules of a CFG. These results together indicate the existence of some LZ-based method to encode both the object and its assembling process. Indeed, the proof of the following Theorem 9 constructs such a LZ compression scheme.
For traditional LZ schemes, the encoded/compressed form of a string comprises a sequence of encoded tuples (or factors) of finite length. Typically, each one of these tuples contains at least one pointer (e.g., in the case of LZ78787878 and LZW, called indexes) that refers to another tuple or to an entry in a dictionary. Also, it can contain information about some elements of the basic alphabet (e.g., in the case of LZ78787878 and LZW, it contains the codeword for only one symbol to be concatenated at the end of the element the pointer is referring to). It is trivial to find examples of objects whose parsing according to the LZ78787878 differs from the parsing according to the LZ77777777, or that the one according to the LZMA differs from LZW, and so on. The LZ family tree of compression schemes includes many distinct schemes, including LZ77, LZ78, LZW, LZSS, and LZMA [62]. All of these compression methods are defined by schemes that resort to a (static or dynamic) dictionary containing tokens, indexes, pointers, or basic symbols from which the decoder or decompressor can retrieve the original raw data from the received encoded/compressed form according to the respective LZ scheme [62]. Theorem 9 demonstrates that the minimum rooted assembly (sub)space from which the assembly index is calculated is a compression scheme that belongs to this LZ family of compression algorithms.
@@
On the one hand, the pointers in traditional LZ compression schemes, such as LZ77, LZ78 or LZW, refer to encoded tuples that have occurred before as recorded by the associated dictionary. On the other hand, for Assembly Theory, the pointers refer to other tuples that might have not occurred before in an assembling path, but are part of the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$, that is, part of a “tree”-like structure that needs to be encoded in the dictionary, which is only built or queried during the encoding/decoding process.
In practice, the major characteristic that distinguishes assembly index compression from other traditional LZ schemes (like LZ77/78 or LZW), is that: in the latter case, one would parse the string into factors, that is, contiguous smaller strings (or codewords) whose resulting unidimensional concatenation directly represents the string; while in the former case, one would parse the object into a non-unidimensional structure where each block is “concatenated” according to a directed acyclic multigraph with edge labels drawn from the set of vertices. For a minimum rooted assembly (sub)space there may be more than one rooted path terminating at the final object to be assembled. Each of these paths corresponds to a distinct way to parse (i.e., decompose) the object.
The sequence of edges that constitutes an assembling path (from the basis set $B_{\Gamma}$ to the terminal object/vertex in a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$) can be converted into a sequence of respective encoded tuples that contains pointers, codewords for other objects, etc. One can think of these assembling paths being, for example, the sequence defined by either one of the longest rooted paths $\gamma_{x}^{\max}$ or one of the shortest rooted paths $\gamma_{x}^{\min}$ in the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$. In particular, the shortest path is often cited by AT’s authors as the one that translates or captures the idea of complexity that the assembly index is aiming at quantifying [1, 41, 53, 2, 8].
Indeed, as demonstrated by the upcoming Theorem 9, the length of the shortest paths is roughly the same as the number of factors (or the number of tuples of pointers) in the encoded form (see Equation (15)) of the LZ scheme that the assembly index calculation method is equivalent to. Thus, as shown in Theorem 9, the LZ compression scheme defined by the assembly index calculation method establishes a way to compress such a non-linear structure (in particular, a directed acyclic multigraph) and non-deterministic parsings/decompositions back into a linear sequence of codewords. We demonstrate that the minimality of ${\Gamma^{*}}_{x}$ in fact guarantees that such a scheme (where pointers may not refer to the unidimensional past subsequences) allows unique decoding, despite the non-unidimensional structure of assembly spaces in which more than one path may lead to the same object.
Theorem 9 (LZ encoding from the assembly index calculation method).
@@
Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly index of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $\Gamma\subset\mathcal{S}$ with $y\in V\left(\Gamma\right)$, $\gamma_{y}^{\min}$, and $\gamma_{y}^{\max}$ be arbitrary. Then, one has it that the sequence $S_{AT}\left(y\right)$ composed of $\left|\gamma_{y}^{\min}\right|$ encoded tuples in the form $c,i_{t},v_{l}\right>$ (except for the 00-th initial tuples which are $t\right>$, $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, and $v_{B}\right>$, respectively, where $v_{B}\in B_{\Gamma}$ and $t\in\mathbb{P}\left(\mathcal{T}\right)$) encodes the assembling process of the object $y$ such that
@@
  • $t\right>$ is an encoding of the $t$-th member of $\mathbb{P}\left(\mathcal{T}\right)$, where $\mathbb{P}\left(\mathcal{T}\right)$ is the powerset of possibles edges of the minimum directed acyclic multigraph $\mathcal{T}$ into which any possible minimum rooted assembly subspace (whose longest rooted paths have length $\left|\gamma_{y}^{\max}\right|$) can be embedded, and all the most distant vertices from the terminal vertex (i.e., $y$) can only be the vertices in $B_{\Gamma}$.
@@
  • $v_{l}$ is another index/pointer that refers to the edge $i^{\prime}_{t}$ whose source vertex $s\left(i^{\prime}_{t}\right)=v\in V\left(\Gamma\right)$ is the vertex that is the label of the edge $i_{t}$ when $\Gamma$ is embedded into $\mathcal{T}$;
@@
  • $c_{\Gamma}\left(y\right)$ gives the assembly index of the object $y$ in the assembly space $\Gamma$ (or $\mathcal{S}$);
@@
as input. Then, by employing $\mathbf{F}$, it constructs the directed acyclic multigraph (DAmG) $\mathcal{T}$ of height $\left|\gamma_{y}^{\max}\right|$, and enumerates all the possible minimum rooted assembly subspaces contained in $\mathcal{T}$ as the following procedure: it starts from those that are isomorphic to an arbitrarily fixed $\gamma_{y}^{\max}$ with any possible combination of initial vertex/object in $B_{\Gamma}$; and then in a dovetailing procedure, it indexes first those subspaces that differ (from this $\gamma_{y}^{\max}$) by the least amount of additional edges. Then, program $p_{5}$ chooses the $t$-th first $\Gamma^{\prime}$ that appears in this enumeration. In the next step, program $p_{5}$ reads the rest of the sequence $S_{AT}\left(y\right)$, extracting from the encoded tuples the sequence $I=\left({i_{t}}_{1},\dots,{i_{t}}_{\left|\gamma_{y}^{\min}\right|},{v_{l}}_{1},\dots,{v_{l}}_{\left|\gamma_{y}^{\min}\right|}\right)$ of indexes of edges that are contained in this $t$-th first $\Gamma^{\prime}$ such that it constitutes a shortest rooted path that starts at vertex $v_{B}$. Finally, it uses this shortest rooted path to assemble the final object $y^{\prime}$ accordingly. This describes the decoding procedure from $S_{AT}\left(y\right)$. In order to construct a program $p^{\prime}_{5}$ that encodes a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ into a $S_{AT}\left(y\right)$, one just needs to note that the codewords for $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, and $y\right>$ are unique and always retrievable by an algorithm given ${\Gamma^{*}}_{y}$ as input. From these, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to construct $\mathcal{T}$. Program $p^{\prime}_{5}$ then employs the same procedures that $p_{5}$ ran in order to enumerate the minimum rooted assembly subspaces embedded into $\mathcal{T}$ so as to find the index $t$ that corresponds to ${\Gamma^{*}}_{y}$ in this enumeration. Next, it enumerates all the possible shortest rooted path in this $t$-th minimum rooted assembly subspace. Then, it extracts $v_{B}$ and $I$ from $\mathcal{T}$ by matching those that correspond to the edges and their respective labels in the first shortest rooted path $\gamma_{y}^{\min}$ found in the previous enumeration that terminates at the object $y$. Finally, the program returns the sequence $S_{AT}\left(y\right)$ as output. ∎
@@
With the purpose of demonstrating that $\Gamma^{\prime}={\Gamma^{*}}_{y}$, first note that there is only a finite number of rooted assembly spaces whose most distant vertices (from the vertex/object $y$) belong to $B_{\Gamma}$ (which is also finite and fixed given $\Gamma$) and are separated by $\left|\gamma_{y}^{\max}\right|$ hops. Therefore, since $\left|\gamma_{y}^{\max}\right|\right>$ and $B_{\Gamma}\right>$ are unique from ${\Gamma^{*}}_{y}$ and the enumeration procedure is the same in both the encoding or the deconding phase, there is a unique index $t$ that corresponds to $\Gamma^{\prime}={\Gamma^{*}}_{y}$. In order to achieve the proof of Equation (14), one can exploit the minimality of ${\Gamma^{*}}_{y}$ and $\gamma_{y}^{\min}$ and the maximality of $\gamma_{y}^{\max}$. From any rooted assembly space one of whose longest rooted paths is $\gamma_{y}^{\max}$, in order to construct $\mathcal{T}$, an algorithm can minimise the number of sufficient and necessary new vertices that needs to be added to $\gamma_{y}^{\max}$ by the following procedure: assign an index $j\in\mathbb{N}$ to each of the edges in $\gamma_{y}^{\max}$ by starting from the closest ones to the farthest ones from the initial vertex, which is the object in the basis set $B_{\Gamma}$, such that $1\leq j\leq\left|\gamma_{y}^{\max}\right|$; for the label of the edge $j=1$ in $\gamma_{y}^{\max}$, one only requires at most one extra vertex, and this vertex can only belong to $B_{\Gamma}$ because otherwise, since there would need to be an additional edge connecting this new vertex to the target vertex of edge $j=1$, it would contradict the maximality of $\gamma_{y}^{\max}$; the same analogously holds for each arbitrary $j+1\leq\left|\gamma_{y}^{\max}\right|$, so that labelling this edge $j+1$ with a vertex only requires at most one extra new vertex, and this vertex can only be at level $\leq j$, i.e., it can only belong to the set of vertices that are separated from the terminal vertex $y$ by at least $\left|\gamma_{y}^{\max}\right|-j+1$ hops (otherwise, it would again contradict the maximality of $\gamma_{y}^{\max}$); and for each new vertex created at level $j$ one would only need at most $2$ new vertices at level $\leq j-1$, and so on. Thus, by induction, one will have it that